OpenStudy (anonymous):

(√3) cosΘ -2sinΘcosΘ=0 answer: 0º, 45º, 135º, 180º but i got 0º, 45º, 135º, 180º, 360º. Why 360º is rejected?

OpenStudy (mertsj):

Did the problem specify a domain? Perhaps they only want answers in this interval: \[0^{o}\le \theta <360^{o}\]

OpenStudy (anonymous):

that was my thought too, but i don't think 360 work

OpenStudy (anonymous):

\(\cos(360)=\cos(0)=1\) and so \(\sqrt{3}\cos(360)=\sqrt{3}\)

OpenStudy (mertsj):

But then 180 won't work for the same reason

OpenStudy (anonymous):

true enough

OpenStudy (anonymous):

\[\cos(x)\left(\sqrt{3}-2\sin(x)\right)=0\] \[\cos(x)=0\] or \[\sin(x)=\frac{\sqrt{3}}{2}\] now that i look, none of these are right

OpenStudy (anonymous):

maybe theses are the answers to a different question

OpenStudy (mertsj):